OpenStudy (anonymous):

PLease help will fan and medal!!!! Write the expression in exponential form. log(3)9 = 2

OpenStudy (anonymous):

\[\log_{3} 9=2\]

OpenStudy (anonymous):

3^2=9

OpenStudy (anonymous):

Exponential form: b^x=y Logarithmic form: logby=x Plug in the numbers. Log form: log3(9)=2 Exp form: 3^2=9

OpenStudy (anonymous):

thank you @mathway

OpenStudy (anonymous):

no problem!